write in standard form the equation of the line that has a slope of 2 and passes through the poing (3,9).
I think you use the y-y1=m(x-x1)

Answer by subudear(62)

You are right you shuld use y - y1 = m(x - x1) to find the line with slope 2 and passing through (3,9)
given m = 2 and x1 = 3, y1 = 9, replacing the values in equation we get
y - 9 = 2(x - 3)
y - 9 = 2x - 6
y = 2x + 3